OpenStudy (anonymous):

Find 4x − 2y if x = −1 and y = −3 My answer is -10 Am I right?

OpenStudy (anonymous):

no

OpenStudy (anonymous):

can you do it step by step so i can show you your mistake

OpenStudy (anonymous):

Is it 2?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

Thank you! sorry it took me a sec to find my mistake

OpenStudy (anonymous):

so heres how you solve this. first do 4 times x, x=-1 so 4x -1= -4. then do 2 times y. y=-3, so 2 x -3= -6. Now do -4 - (-6)=2